# GLOBALIZATION? TRADE WAR? A COUNTERBALANCE PERSPECTIVE

### v4

@@ 2.4 The Matthew Effect in $\pi$
In Eq. (4), country $i$ accrues its competitiveness in proportion to its initial level of competitiveness, i.e.,
$\pi_{i}=\pi_{i}P_{ii}+\sum\limits_{j\not=i}\pi_{j}P_{ji},$ (7)
@@
in which the proportion $P_{ii}$ is generally larger than .7.7.7.7. Thus, $\pi$ has the Matthew effect of accumulated advantage, sometimes summarized by the aphorism “the rich get richer and the poor get poorer.” Consequently, powerful economies have substantial competitiveness compared to less powerful ones, revealing a nonlinear relation between $\pi_{i}$ and $i$’s production $g_{i}$. For example, the USA’s $\pi_{i}$ could be much higher than Japan’s, compared to their production ratio. The amplification also comes from the USA’s relative advantages over other countries, compared to Japan’s. In Eq. (7) with $i=\mathrm{USA}$ or $i=\mathrm{Japan}$,
@@
$\left\{\begin{array}[]{lcl}\pi_{i}^{\mathrm{USA}}&=&\frac{1}{1-P_{ii}^{\mathrm{USA}}}\sum\limits_{j\not=\mathrm{USA}}\pi_{j}P_{ji}^{\mathrm{USA}},\\ \pi_{i}^{\mathrm{Japan}}&=&\frac{1}{1-P_{ii}^{\mathrm{Japan}}}\sum\limits_{j\not=\mathrm{Japan}}\pi_{j}P_{ji}^{\mathrm{Japan}},\end{array}\right.$
@@
the Mathew effect in $\pi_{i}^{\mathrm{USA}}$ comes from two inequalities: $01>P_{ii}^{\mathrm{USA}}>P_{ii}^{\mathrm{Japan}}>$ 0; $P_{ji}^{\mathrm{USA}}>P_{ji}^{\mathrm{Japan}}>0$, in general, for other countries $j$. Alternatively, the Matthew effect is largely explained by the preferential attachment in the endogenous weighting, whereby the weight is proportionally distributed among the countries according to how competitive they already are.
@@
The effect implies that $\pi_{i}/\pi_{j}>g_{i}/g_{j}$ when $g_{i}$ is significantly larger than $g_{j}$. If we normalize $g_{i}$ by $\tilde{g}_{i}\overset{\mathrm{def}}{=\joinrel=}g_{i}/\sum\limits_{j=1}^{n}g_{j}$, then the effect also means that $\pi_{i}$ positively correlates with $\tilde{g}_{i}^{2}$ after controlling $\tilde{g}_{i}$. So, it has a quadratic growth when $\tilde{g}_{i}$ soars linearly and in a linear regression,
$\pi_{i}=c_{1}\tilde{g}_{i}+c_{2}\tilde{g}_{i}^{2}+\epsilon_{i}$ (8)
@@
for some positive $c_{1}$ and $c_{2}$. Therefore, when dealing with a rapidly thriving country $j$, $i$ should mitigate its conflicts with $j$ before the counterpart becomes too strong to compromise. Furthermore, if the country capitalizes on the accelerating $\pi_{i}$ in its production $g_{i}$, say,
$g_{i}=c_{3}+c_{4}\pi_{i}+\epsilon_{i},\quad\mathrm{for}\ \mathrm{some}\ c_{4}>0,$
@@
then $g_{i}$ would be accelerated quadratically, giving rise to $\pi_{i}$’s quartic growth. Continuing this iteration, eventually, both $\pi_{i}$ and $g_{i}$ would grow exponentially to their maximum capacity. Many empires have suddenly emerged from the horizon with speedily burgeoning importance in human history.
@@
Granting extra welfare to competitive economies is another major force that drives trade flows over the network. So, an economic superpower would make all efforts to maintain its superpower status when facing challenges from emerging competitors; maintaining its advantage in $\pi$ avoids the domino effect, the opposite of the Matthew effect. As Copeland (2014) says, “When expectations [of the future trade environment] turn negative, leaders are likely to fear a loss of access to raw materials and markets, giving them an incentive to initiate crises to protect their commercial interests.” Thus, their incentive to start crises is more proportional to $\pi_{i}$ than $g_{i}$. Accordingly, an emerging superpower’s aspiration to assume its rightful place as a superpower would face serious challenges. World War I and II were fiery combats between existing and emerging superpowers, resetting new world orders.
The welfare of competitive advantage leads to an agglomeration effect, which clusters all countries around a handful of economic superpowers. For example, country $i$ considers certain imports from either country $j$ or $k$ where $g_{j}$ is significantly larger than $g_{k}$, other things being equal. If $i$ selects $j$, by Eq. (4), then the competitiveness country $i$ derives from $j$ in the imports is approximately
Because of the Matthew effect, ${\pi_{j}}/{g_{j}}>{\pi_{k}}/{g_{k}}$ and, so, country $i$ would likely choose $j$, other things being equal. Otherwise, $k$ could counter the offer by lowering its price of exports to $(\pi_{k}g_{j})/(\pi_{j}g_{k})$ of $j$’s price. In either case, a small country (such as $k$) tends to fear the emergence of a neighboring superpower (such as $j$) that shares similar location advantage and resource endowment.
Finally, Eq. (3) implies that $\pi$ is the unit-sum row eigenvector of the largest row eigenvalue of $P$. It is a type of eigenvector centrality, a subclass of network centrality. In the literature, eigenvector centrality is the dominant eigenvector for an adjacency or relation matrix, which often has a zero diagonal (e.g., Bonacich, 1987). Our $P$ has a nonzero diagonal such that domestic and foreign goods can compete for the domestic market, and the eigenvector $\pi$ measures the competitiveness. The same domestic products are used domestically or overseas. Hu, (2020) also retains a nonzero diagonal when any accepted student selects only one college to attend from multiple offers; so, the selected college on the diagonal competes with the others that also accept the student and are off the diagonal. Besides, $P$ normalizes the scales of economies such that each row has a unit sum. As a result, $P$ has the largest row eigenvalue $1$, which is common to all stochastic matrices, and its unit-sum eigenvector $\pi$, which is specific to $P$.
@@ 3 Bilateral Trade War for Competitiveness
Expecting a positive change to $\pi_{i}$, country $i$ considers a trade war against its trade partner $j$. Gaming on the matrix $P$, $i$’s first-move action could change the element $P_{ji}$ — the portion of $j$’s production exported to $i$. Raising tariffs or decreasing import quotas upon $j$, for example, reduces imports from country $j$ and decreases $P_{ji}$. Country $i$ could also restrict certain exports to $j$, directly decreasing $P_{ij}$. For simplicity, we assume that the bilateral conflict involves only the two counterparties without directly sacrificing or benefiting $P_{ik},P_{ki},P_{jk}$, or $P_{kj}$ for any third party $k$. Nevertheless, the change on $P_{ji}$ or $P_{ij}$ eventually ramifies over the global value chains in multiple steps, and all third parties get involved indirectly. The simplicity could be far from realistic in the 2018 China-USA trade war as the Biden administration also urged its allies to ban Chinese companies from buying advanced chips and chip-making equipment (e.g., New York Times, 2022).
@@
As a reprisal, country $j$ could react by decreasing $P_{ij}$ or $P_{ji}$ sooner or later. Without the reprisal, the war dies out at inception, and country $i$ wins. No reaction or unconditional surrender could avoid further damages, but it could also bring in further sanctions from $i$ when $j$ shows a sign of weakness. Overreaction to the hostile change on $P_{ji}$ or $P_{ij}$ may also severely hurt $\pi_{j}$. After the Trump administration began setting tariffs and other trade barriers on China in January 2018, for example, China responded in April 2018 by imposing tariffs on 128 products imported from America (cf., Washington Post, 2018). Say, in retaliation for a $\Delta P_{ji}$ change on $P_{ji}$, country $j$ changes $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$, i.e., $\Delta P_{ij}=\lambda_{ji}\Delta P_{ji}$. Similarly, in retaliation for $i$’s $\Delta P_{ij}$ change on $P_{ij}$, $j$ changes $P_{ji}$ by $\lambda_{ij}\Delta P_{ij}$, i.e., $\Delta P_{ji}=\lambda_{ij}\Delta P_{ij}$. So, $\lambda_{ij}=1/\lambda_{ji}$ in the veil of any first-mover advantage. We assume that $\lambda_{ji}>0$ to align the actions from both countries. Thus, ignoring the temporally lagged reactions in the long-run effects on $\pi$, changing $P_{ji}$ by $\Delta P_{ji}$ and changing $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$ have the same impact on $\pi$. Finally, to maintain the unit sum in each row of $P$, we need to deduct $P_{jj}$ by $\Delta P_{ji}$ and $P_{ii}$ by $\Delta P_{ij}$; so the reduced exports will be consumed domestically, and the modified $P$ remains as a stochastic matrix. We let $\lambda_{ii}=1$ and place all $\lambda_{ij}$ into a $n\times n$ matrix $\Lambda=[\lambda_{ij}]$, $\forall i,j\in\mathcal{N}$.
@@
The coefficient $\lambda_{ji}$ often depends on many exogenous factors from both sides, such as political considerations, anti-dumping protection, trade deficits, and specific trade commodities or services. In general, there is no specific formula for every contingency, and negotiations in practice could involve many rounds of bilateral talks. For a specific context with a set of fair assumptions, however, many researchers in the literature have studied bilateral or even multilateral negotiations (e.g., Bagwell, Staiger, and Yurukoglu, 2020; Schneider, 2005). In the following exposition, we first assume the coefficient is an exogenous constant in the potential trade war. Then, Section 4 studies a generic but simple bargaining solution of $\lambda_{ji}$ for most of the $n(n-1)/2$ conflicts. The solution aims to mitigate the conflicts and fairly share the costs rather than eliminate them; it serves the diverse interests of the countries. In the mixed cooperative and noncooperative context, simply maximizing $\lambda_{ji}$ or $\lambda_{ij}$ could be of little interest to both counterparts.
@@
With the above setting and notations, for any $j\not=i$,
@@
Capitalizing on the relation $\mathrm{d}P_{ij}=\lambda_{ji}\mathrm{d}P_{ji}$ and $\lambda_{ij}={1}/{\lambda_{ji}}$, we also obtain
@@
Using Theorem 2, we have the following Corollary 1, already implicated in Eq. (6) —  $\pi_{i}$ remains unchanged when changed authority flow from $i$ to $j$ equals that from $j$ to $i$. However, Eq. (6) applies no assumptions in Theorem 2.
@@
has all non-negative elements, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ and $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$ have opposite signs, as stated in Corollary 2. In particular, when $\lambda_{ji}=0$ in Eq. (9)-(10), dumping goods and services to country $j$ without resistance would improve $\pi_{i}$ and deteriorate $\pi_{j}$. Also, Eq. (10) may imply that $\lambda_{ji}$ measures $j$’s negotiation power with $i$ in the trade friction if $\pi_{j}$ is $j$’s sole objective.
@@
For any $i\not=j$, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}\leq 0$ and $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}\leq 0.$
By varying $j$ in $\mathcal{N}$, we could apply Eq. (9) to identify country $i$’s trade collaborators and competitors. In Eq. (9), $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ has an opposite sign of $\lambda_{ji}\pi_{i}-\pi_{j}$. If $\lambda_{ji}>\pi_{j}/\pi_{i}$, then $j$ is a potential target for conflict with $i$ because a negative infinitesimal $\mathrm{d}P_{ji}$ or a negative small $\Delta P_{ji}$ would bring a positive $\mathrm{d}\pi_{i}$ or $\Delta\pi_{i}$, respectively. Similarly, the condition $\lambda_{ji}<\pi_{j}/\pi_{i}$ implies that $j$ is a potential target for further collaboration. Whenever $\lambda_{ji}<\pi_{j}/\pi_{i}$ and $\mathrm{d}P_{ji}>0$, the larger $\lambda_{ji}$, the slower $\pi_{i}$ grows and $\pi_{j}$ shrinks. When $\lambda_{ji}>\pi_{j}/\pi_{i}$, $i$ could also increase $\pi_{i}$ by adjusting the sign of $\mathrm{d}P_{ji}$. If $j$ insists on $\lambda_{ji}=\pi_{j}/\pi_{i}$, then any tiny change on $P_{ji}$ would have no significant effects on both $\pi_{i}$ and $\pi_{j}$ regardless of the sign of $\mathrm{d}P_{ji}$.
In search for the best competitor and collaborator, we make the results comparable across countries by taking percentage changes in both $\pi_{i}$ and $P_{ji}$. The limit of the percentage change of $\pi_{i}$ with respect to that of $P_{ji}$ is
@@
$\mathop{\mathrm{argmax}}\limits_{j\not=i}\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}$ (13)
as $i$’s best trade partner and
$\mathop{\mathrm{argmin}}\limits_{j\not=i}\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}$ (14)
@@
as $i$’s best competitor. In Eq. (13) or (14), $\pi_{i}$ has the largest increase or decrease, respectively, in percentage per one percentage rise of $P_{ji}$. However, $i$ is not his best partner’s best partner nor his best competitor’s best competitor, according to Corollary 2. Thus, $i$ and $j$ are not in a sports-like competition or zero-sum game unless they are evenly matched. For example, in the 2018 China-USA trade war, both the Trump and Biden administrations actively challenged China, whereas China passively defended itself.
@@
Eq. (11) is the side effects on all other countries due to $i$’s change on $P_{ji}$, including $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$. With the perturbation on $P_{ji}$, some countries gain, and others lose their competitiveness; the gain or loss may be even more substantial than $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$. Though $\pi_{i}$ and $\pi_{j}$ always move in opposite directions, the movements are not the same magnitude, so one movement could be significant, whereas the other remains insignificant. Third parties unevenly share the difference between these two movements, and their individual or aggregate shares may be significant.
@@
The side effects may play a role in determining $\lambda_{ji}$ when $i$ is much more potent than $j$. The conflict seems resolved if the first-order condition $\lambda_{ji}=\pi_{j}/\pi_{i}$ maximizes $\pi_{i}$. However, if $\lambda_{ji}=\pi_{j}/\pi_{i}$ maximizes $\pi_{j}$ instead, then in theory, $i$ could extremely counteract and push $\lambda_{ji}$ near zero. In practice, however, $i$ unlikely exercises these extreme actions because $\pi_{i}$’s gain from $j$ is capped by $\pi_{j}P_{ji}$, as seen from Eq. (4), which could be negligible to $i$. Also, $\pi_{j}P_{ji}$ could be less than the loss of third parties, measured by Eq. (11). So, excessively non-cooperative counteractions could provoke outrage from these countries, and $j$ could ally with them to seek collective bargaining with $i$. It could also pursue an objective other than maximizing $\pi_{j}$, avoiding confrontation with $i$. For example, it can engage in country $i$’s economy by developing complementary industries and focusing on its comparative advantages. For centuries, many norms and standards have been set up to tame trade bullies; today, the World Trade Organization (WTO) deals with trade rules between nations.
4 A Nash Bargaining Solution for Reprisal Coefficients $\Lambda$
When other factors also play a role, the condition $\lambda_{ji}>\pi_{j}/\pi_{i}$ alone is insufficient for country $i$ to start a conflict action against $j$. Among them, the most primary are trade deficits. This section studies a bilateral bargaining solution for $\lambda_{ji}$ when trade balance is also an objective. If small economies are willing to dodge direct competition with large ones, the vast majority of the $n(n-1)/2$ conflicts are resolved or mitigated by the bargaining solution.

### v5

@@ 2.4 The Matthew Effect in $\pi$
In Eq. (4), country $i$ accrues its competitiveness in proportion to its initial level of competitiveness:
$\pi_{i}=\pi_{i}P_{ii}+\sum\limits_{j\neq i}\pi_{j}P_{ji},$ (7)
@@
where the proportion $P_{ii}$ is generally larger than .7.7.7.7. This illustrates $\pi$’s Matthew effect of accumulated advantage, often summarized by the aphorism “the rich get richer and the poor get poorer.” Consequently, powerful economies exhibit substantial competitiveness compared to less powerful ones, revealing a quadratic relationship between $\pi_{i}$ and $g_{i}$. For instance, the USA’s $\pi_{i}$ could be significantly higher than Japan’s, relative to their production ratio. The amplification also stems from the USA’s relative advantages over other countries compared to Japan’s. In Eq. (7), with $i=\mathrm{USA}$ or $i=\mathrm{Japan}$, respectively,
@@
$\left\{\begin{array}[]{lcl}\pi_{i}^{\mathrm{USA}}&=&\frac{1}{1-P_{ii}^{\mathrm{USA}}}\sum\limits_{j\neq\mathrm{USA}}\pi_{j}P_{ji}^{\mathrm{USA}},\\ \pi_{i}^{\mathrm{Japan}}&=&\frac{1}{1-P_{ii}^{\mathrm{Japan}}}\sum\limits_{j\neq\mathrm{Japan}}\pi_{j}P_{ji}^{\mathrm{Japan}},\end{array}\right.$
@@
the Mathew effect in $\pi_{i}^{\mathrm{USA}}$ arises from two inequalities: $01>P_{ii}^{\mathrm{USA}}>P_{ii}^{\mathrm{Japan}}>$ 0; $P_{ji}^{\mathrm{USA}}>P_{ji}^{\mathrm{Japan}}>0$, generally for other countries $j$. Alternatively, the Matthew effect is largely explained by preferential attachment in the endogenous weighting, whereby the weight is proportionally distributed among countries according to their existing competitiveness.
@@
The effect implies that $\pi_{i}/\pi_{j}>g_{i}/g_{j}$ when $g_{i}$ is significantly larger than $g_{j}$. If we normalize $g_{i}$ by $\tilde{g}_{i}\overset{\mathrm{def}}{=\joinrel=}g_{i}/\sum\limits_{j=1}^{n}g_{j}$, then the effect also means that $\pi_{i}$ positively correlates with $\tilde{g}_{i}^{2}$ after controlling $\tilde{g}_{i}$. Thus, it exhibits quadratic growth when $\tilde{g}_{i}$ increases linearly, and in a linear regression:
$\pi_{i}=c_{1}\tilde{g}_{i}+c_{2}\tilde{g}_{i}^{2}+\epsilon_{i},$ (8)
@@
for some $c_{1}>0,c_{2}>0$, where $\epsilon_{i}$ is the residual term. Therefore, when dealing with a rapidly thriving country $i$, country $j$ should mitigate its conflicts with $i$ before the counterpart becomes too strong to compromise. Furthermore, if the country capitalizes on the accelerating $\pi_{i}$ in its production $g_{i}$, say:
$\tilde{g}_{i}=c_{3}+c_{4}\pi_{i}+\epsilon_{i}$
@@
for some $c_{4}>0$, then $\tilde{g}_{i}$ would be accelerated quadratically, leading to $\pi_{i}$’s quartic growth. Continuing this iteration, eventually, both $\pi_{i}$ and $\tilde{g}_{i}$ would grow exponentially to their maximum capacities. Many empires have suddenly emerged from the horizon with speedily burgeoning importance in human history.
Additionally, the welfare associated with competitive advantage leads to an agglomeration effect, clustering all countries around a select few of economic superpowers. According to Corollary 1, imports from a large economy are preferred over imports from a smaller economy, all else being equal. Consequently, the smaller economy should counter this preference by lowering its export prices. Meanwhile, a third party could impose different tariffs on large and small countries to ensure that the same imported products have uniform prices in the third country. Furthermore, the large market, characterized by numerous producers and consumers where no single entity can dominate, may be favored by the third party for its market competitiveness. In either scenario, a small country tends to fear the emergence of a neighboring superpower that shares similar location advantages and resource endowments.
Corollary 1
Assume the Matthew effect. If $g_{j}>g_{k}$, then the third party $i$ prefers imports from country $j$ over country $k$, all else being equal.
@@
Thus, providing additional welfare to competitive economies is another significant force driving trade flows across the network, alongside comparative advantage. However, trade wars and protectionism are more likely due to competitive advantages rather than comparative advantages, which benefit all trade partners. Accordingly, an economic superpower would exert all efforts to maintain its status when facing challenges from emerging competitors; preserving its advantage in $\pi$ prevents the domino effect, which is contrary to the Matthew effect. As Copeland (2014) states, “When expectations [of the future trade environment] turn negative, leaders are likely to fear a loss of access to raw materials and markets, giving them an incentive to initiate crises to protect their commercial interests.” Hence, their incentive to start crises is more proportional to $\pi_{i}$ than $g_{i}$. An emerging superpower’s aspiration to assume its rightful place would encounter significant challenges. World War I and II were intense conflicts between existing and emerging superpowers, leading to the establishment of new world orders.
@@ 3 Bilateral Trade War for Competitiveness
Anticipating a significant positive change to $\pi_{i}$, country $i$ considers initiating a trade war against its trade partner $j$. Gaming on the matrix $P$, country $i$’s first move could alter the element $P_{ji}$ — the portion of $j$’s production exported to $i$. For instance, raising tariffs or decreasing import quotas on $j$ reduces imports from $j$ and decreases $P_{ji}$. Country $i$ could also restrict certain exports to $j$, directly decreasing $P_{ij}$. For simplicity, we assume the bilateral conflict involves only the two counterparties in $P$ without directly affecting $P_{ik},P_{ki},P_{jk}$, or $P_{kj}$ for any third party $k$. However, changes in $P_{ji}$ or $P_{ij}$ eventually impact global value chains in multiple steps, indirectly involving all third parties. This simplification may not reflect the complexity of the 2018 China-USA trade war, where the Biden administration urged allies to ban Chinese companies from purchasing advanced chips and chip-making equipment (New York Times, 2022).
In retaliation, country $j$ might decrease $P_{ij}$ or $P_{ji}$ sooner or later. Without retaliation, the war ends at inception, and country $i$ wins. No reaction or unconditional surrender could avoid further damages but might also invite additional sanctions from $i$ if $j$ shows weakness. Overreaction to hostile changes in $P_{ji}$ or $P_{ij}$ may also severely harm $\pi_{j}$. For example, after the Trump administration imposed tariffs and trade barriers on China in January 2018, China responded in April 2018 by imposing tariffs on 128 products imported from the USA (Washington Post, 2018).
@@
We introduce a matrix $\Lambda$ for retaliation actions. In retaliation for a change $\Delta P_{ji}$ in $P_{ji}$, country $j$ changes $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$, i.e., $\Delta P_{ij}=\lambda_{ji}\Delta P_{ji}$. Similarly, in retaliation for $i$’s change $\Delta P_{ij}$ in $P_{ij}$, $j$ changes $P_{ji}$ by $\lambda_{ij}\Delta P_{ij}$, i.e., $\Delta P_{ji}=\lambda_{ij}\Delta P_{ij}$. Thus, $\lambda_{ij}=1/\lambda_{ji}$, in the veil of any first-mover advantage. We assume $\lambda_{ji}>0$ to align actions from both countries. Ignoring temporally lagged reactions in the long-run effects on $\pi$, changing $P_{ji}$ by $\Delta P_{ji}$ and changing $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$ have the same impact. We let $\lambda_{ii}=1$ and place all $\lambda_{ij}$ into an $n\times n$ matrix $\Lambda=[\lambda_{ij}]$, $\forall i,j\in\mathcal{N}$. Finally, to maintain the unit sum in each row of $P$, we need to deduct $P_{jj}$ by $\Delta P_{ji}$ and $P_{ii}$ by $\Delta P_{ij}$; thus, the reduced exports will be consumed domestically, and the modified $P$ remains a stochastic matrix.
@@
The coefficient $\lambda_{ji}$ often depends on various exogenous factors from both sides, such as political considerations, anti-dumping protection, trade deficits, and specific trade commodities or services. Generally, there is no specific formula for every contingency, and negotiations in practice could involve many rounds of bilateral talks. In the following exposition, we first assume the coefficient is an exogenous or predetermined constant in the potential trade war. Then, Section 4 studies a generic but simple bargaining solution of $\lambda_{ji}$ for most of the $n(n-1)/2$ conflicts. The solution aims to mitigate the conflicts and fairly share the costs rather than eliminate them; it serves the diverse interests of the countries. In the mixed cooperative and noncooperative context, simply maximizing $\lambda_{ji}$ or $\lambda_{ij}$ could be of little interest to both counterparts.
@@
With the above setting and notations, for any $j\neq i$,
@@
Capitalizing on the relation $\mathrm{d}P_{ij}=\lambda_{ji}\mathrm{d}P_{ji}$ and $\lambda_{ij}={1}/{\lambda_{ji}}$, we also obtain:
@@
Using Theorem 2, we have the following Corollary 2, already implied in Eq. (6) —  $\pi_{i}$ remains unchanged when changed authority flow from $i$ to $j$ equals that from $j$ to $i$. However, Eq. (6) applies no assumptions in Theorem 2.
@@
has all non-negative elements, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ and $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$ have opposite signs, as stated in Corollary 3. In particular, when $\lambda_{ji}=0$ in Eqs. (9)-(10), dumping goods and services to country $j$ without resistance would improve $\pi_{i}$ and deteriorate $\pi_{j}$. Eq. (10) may imply that $\lambda_{ji}$ measures $j$’s negotiation power with $i$ in the trade friction if $\pi_{j}$ is $j$’s sole objective.
@@
For any $i\neq j$, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}\leq 0$ and $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}\leq 0.$
By varying $j$ in Eq. (9), we can identify country $i$’s trade collaborators and competitors. In Eq. (9), $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ has an opposite sign of $\lambda_{ji}\pi_{i}-\pi_{j}$. If $\lambda_{ji}>\pi_{j}/\pi_{i}$, then $j$ is a potential target for conflict with $i$ because a negative infinitesimal $\mathrm{d}P_{ji}$ or a negative small $\Delta P_{ji}$ would bring a positive $\mathrm{d}\pi_{i}$ or $\Delta\pi_{i}$, respectively. Similarly, the condition $\lambda_{ji}<\pi_{j}/\pi_{i}$ implies that $j$ is a potential target for further collaboration. Whenever $\lambda_{ji}<\pi_{j}/\pi_{i}$ and $\mathrm{d}P_{ji}>0$, the larger $\lambda_{ji}$, the slower $\pi_{i}$ grows and $\pi_{j}$ shrinks. When $\lambda_{ji}>\pi_{j}/\pi_{i}$, $i$ could also increase $\pi_{i}$ by adjusting the sign of $\mathrm{d}P_{ji}$. If $j$ insists on $\lambda_{ji}=\pi_{j}/\pi_{i}$, then any tiny change in $P_{ji}$ would have no significant effects on both $\pi_{i}$ and $\pi_{j}$, regardless of the sign of $\mathrm{d}P_{ji}$.
In search of the best competitor and collaborator, we make the results comparable across countries by taking percentage changes in both $\pi_{i}$ and $P_{ji}$. The limit of the percentage change of $\pi_{i}$ with respect to that of $P_{ji}$ is
@@
Furthermore, country $i$ would choose:
$\mathop{\mathrm{argmax}}\limits_{j\neq i}\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}$ (13)
as $i$’s best trade partner and:
$\mathop{\mathrm{argmin}}\limits_{j\neq i}\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}$ (14)
@@
as $i$’s worst competitor. In Eqs. (13) or (14), $\pi_{i}$ has the largest increase or decrease, respectively, in percentage per one percentage rise of $P_{ji}$. However, $i$ is not his best partner’s best partner nor his worst competitor’s worst competitor, according to Corollary 3. Thus, countries $i$ and $j$ are not in a sports-like competition or zero-sum game unless they are evenly matched. For example, in the 2018 China-USA trade war, both the Trump and Biden administrations actively challenged China, whereas China passively defended itself.
@@
Eq. (11) describes the side effects on all other countries due to country $i$’s change in $P_{ji}$. With the perturbation in $P_{ji}$, some countries gain, while others lose their competitiveness; the gain or loss may be more substantial than $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$. Although $\pi_{i}$ and $\pi_{j}$ always move in opposite directions, the movements are not of the same magnitude. One movement could be significant, while the other remains insignificant; it is also possible that both movements are insignificant. Third parties unevenly share the difference between these two movements, and their individual or aggregate shares may be significant.
@@
The side effects may play a role in determining $\lambda_{ji}$ when $i$ is much more potent than $j$. The conflict seems resolved if the first-order condition $\lambda_{ji}=\pi_{j}/\pi_{i}$ maximizes $\pi_{i}$. However, if $\lambda_{ji}=\pi_{j}/\pi_{i}$ maximizes $\pi_{j}$ instead, then in theory, $i$ could extremely counteract and push $\lambda_{ji}$ near zero. In practice, however, $i$ is unlikely to exercise these extreme actions because its gain from $j$ is capped by $\pi_{j}P_{ji}$, as seen from Eq. (4), which could be negligible to $i$. Also, $\pi_{j}P_{ji}$ could be less than the loss of third parties, measured by Eq. (11). So, excessively non-cooperative counteractions could provoke outrage from these countries, and $j$ could ally with them to seek collective bargaining with $i$. It could also pursue an objective other than maximizing $\pi_{j}$, avoiding confrontation with $i$. For example, it can engage in country $i$’s economy by developing complementary industries and focusing on its comparative advantages. For centuries, many norms and standards have been established to tame unilateral trade bullies; today, the World Trade Organization (WTO) deals with trade rules between nations.
4 Nash Bargaining Solutions for Reprisal Coefficients $\Lambda$
When additional factors are considered, the condition $\lambda_{ji}>\pi_{j}/\pi_{i}$ alone is insufficient for country $i$ to initiate a conflict against country $j$. Among these factors, trade deficits are paramount. This section explores bilateral bargaining solutions for $\lambda_{ji}$ when trade balance is also an objective.